OpenStudy (dancergirl45):

Which of the following exponential functions goes through the points (1, 6) and (2, 12)? 3(2)^x 2(3)^x 3(2)^-x 2(3)^-x

OpenStudy (dancergirl45):

@jango_IN_DTOWN

OpenStudy (jango_in_dtown):

do you have a graphing calculator?

OpenStudy (jango_in_dtown):

ok wait.. lets do it in four parts

OpenStudy (jango_in_dtown):

1st case Let y=3.2^x when x=1, y= 3.2^1=3.2=6 so (1,6) is a point on it when x=2, y=3.2^2=3.4=12 so(2,12) is a point on it so 3.2^x contains the points (1,6) and (2,12)

OpenStudy (dancergirl45):

So A

OpenStudy (jango_in_dtown):

there may be more than 1 possibilities right? so we have to check for all of them

OpenStudy (jango_in_dtown):

now for the 2nd case y=2.3^x put x=1 and you have y=6 put x=2 and you have y=18 so (2,12) doesnot lie on it. so 2nd case is not possible

OpenStudy (jango_in_dtown):

3rd and 4th case will also be not true.. so only 1st case is true

OpenStudy (dancergirl45):

Thank you so much

OpenStudy (dancergirl45):

hey could u check this

OpenStudy (dancergirl45):

Two different plants grow each year at different rates, which are represented by the functions f(x) = 4x and g(x) = 5x + 2. What is the first year the f(x) height is greater than the g(x) height?

OpenStudy (dancergirl45):

the choices are year 0 year 1 year 2 year 3

OpenStudy (dancergirl45):

i got year 1

OpenStudy (dancergirl45):

is that right

OpenStudy (jango_in_dtown):

may be you wrote f(x) or g(x) wrong

OpenStudy (jango_in_dtown):

5x>4x so 5x+2>4x.. g(x)>f(x) alwaysssssssss

OpenStudy (jango_in_dtown):

@dancergirl45

OpenStudy (dancergirl45):

yes

OpenStudy (jango_in_dtown):

check the functions. they are not correct

OpenStudy (dancergirl45):

i copied and pasted it

OpenStudy (jango_in_dtown):

question is wrong and i justified you the reason. it is never possible that f(x)>g(x)

OpenStudy (dancergirl45):

oh ok